THE HIGH RESOLUTION FTIR SPECTRUM OF THE ν_{3} FUNDAMENTAL OF THE CIS CONFORMER OF METHYL NITRITE

\underline{LISA M. GOSS} , Department of Chemistry, Idaho State University, Campus Box 8023, Pocatello ID, 83209; THOMAS A. BLAKE, Pacific Northwest National Laboratory, P. O. Box 999, Mail Stop K8-88, Richland, WA 99352.

Methyl nitrite (CH_{3}ONO) contains a mixture of cis and trans rotational isomers. As in the ν_{8} and ν_{9} vibrations of the cis form, the ν_{3} vibration is not too congested, consistent with the high barrier to rotation of the methyl group in the cis form. The ν_{3}, with an origin at 1620. cm ^{−1}, has been assigned by previous workers to the N=O stretch of the cis conformer and hot bands are also present in this region. In this study, a spectrum taken at 0.0015 cm ^{−1} resolution on the PNNL IFS120/HR in a 20 cm cell was used to assign more than 300 A-type transitions of the cis ν_{3} between 1600 and 1650 cm ^{−1} and a set of excited state constants was determined.